# Non-stabilizerness generation in a multi-particle quantum walk

### v1

@@ Acknowledgements.
This work received financial support from CNCS/CCCDI–UEFISCDI, under projects number PN-IV-P1-PCE-2023-0159, PN-IV-P1-PCE-2023-0987, PN-IV-P8-8.3-PM-RO-FR-2024-0059, and by the “Nucleu” Program within the PNCDI 2022–2027, Romania, carried out with the support of MEC, project no. 27N/03.01.2023, component project code PN 23 24 01 04. This research was also supported by the National Research, Development and Innovation Office - NKFIH within the Quantum Technology National Excellence Program (Project No. 2017-1.2.1-NKP-2017-00001), K134437, K142179 by the BME-Nanotechnology FIKP grant (BME FIKP-NAT), the QuantERA ‘QuSiED’ grant No. 10101773.

### v2

@@ Acknowledgements.
This work received financial support from CNCS/CCCDI-UEFISCDI, under projects No. PN-IV-P1-PCE-2023-0159, PN-IV-P1-PCE-2023-0987, PN-IV-P8-8.3-PM-RO-FR-2024-0059, and by the “Nucleu” Program within the PNCDI 2022-2027, Romania, carried out with the support of MEC, project no. 27N/03.01.2023, component project code PN 23 24 01 04. This research was also supported by the National Research, Development and Innovation Office - NKFIH within the Quantum Technology National Excellence Program (Project No. 2017-1.2.1-NKP-2017-00001 and grant No. SNN139581), K134437, K142179 by the BME-Nanotechnology FIKP grant (BME FIKP-NAT), and the QuantERA ‘QuSiED’ grant No. 10101773. This work was also supported by the HUN-REN Hungarian Research Network through the Supported Research Groups Programme, HUN-REN-BME-BCE Quantum Technology Research Group (TKCS-2024/34).